OpenStudy (anonymous):

f(x)=cos(x)/(2+sin(x)) find at which points it has a horizontal tangent line on the interval [0,2pi]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

oops i mixed two problems

OpenStudy (anonymous):

\[y'=\frac{-\sin x(2+\sin x)-\cos x(\cos x)}{(2+\sin x)^2}\]simplify

OpenStudy (anonymous):

yeah got i now i mixed this one with the problem before it. thank you!

OpenStudy (anonymous):

then set y'=0 and solve on the interval [0,2pi]

OpenStudy (anonymous):

no problem